# AppReferences

### v2

@@ Abstract
Principal stratification provides a causal inference framework for investigating treatment effects in the presence of a post-treatment variable. Principal strata play a key role in characterizing the treatment effect by identifying groups of units with the same or similar values for the potential post-treatment variable under both treatment levels. The literature has focused mainly on binary post-treatment variables. Few papers considered continuous post-treatment variables. In the presence of a continuous post-treatment, a challenge is how to identify and characterize meaningful coarsening of the latent principal strata that lead to interpretable principal causal effects. This paper introduces the confounders-aware shared-atom Bayesian mixture, a novel approach for principal stratification with binary treatment and continuous post-treatment variables. Our method leverages Bayesian nonparametric priors with an innovative hierarchical structure for the potential post-treatment variable that overcomes some of the limitations of previous works. Specifically, the novel features of our method allow for (i) identifying coarsened principal strata through a data-adaptive approach and (ii) providing a comprehensive quantification of the uncertainty surrounding stratum membership. Through Monte Carlo simulations, we show that the proposed methodology performs better than existing methods in characterizing the principal strata and estimating principal effects of the treatment. Finally, our proposed model is applied to a case study in which we estimate the causal effects of US national air quality regulations on pollution levels and health outcomes.
keywords:
Bayesian Causal Inference; Principal Stratification; Shared Atom Mixture Model.
@@ 1 Introduction
Principal stratification is a well-established causal inference approach that accounts for post-treatment variables (Frangakis & Rubin, 2002). A post-treatment variable is defined as a variable that is potentially affected by treatment and it is also associated with the primary potential outcomes. Thus, a post-treatment variable can be seen as an intermediate variable between the treatment and the outcome of interest. Post-treatment variables may pose inferential challenges for the estimation of treatment effects, but also provide a useful approach to conceptualizing the mechanisms/pathways through which the treatment affects the outcome through the intermediate variable.
@@
Principal stratification with respect to the post-treatment variable, involves stratifying units into different groups, known as principal strata. Principal strata consist of units that share the same or similar potential values of the post-treatment variable under all possible treatment levels: by definition, principal strata are not affected by the treatment and thus represent an inherent latent characteristic of the units.
@@
Since the assignment of treatment does not influence the principal strata, comparisons of potential outcomes at different treatment levels within a principal stratum represent well-defined causal effects (Rubin, 1974). These specific comparisons are referred to as principal causal effects. Researchers have often referred to the dissociative (and associative) effect as the effect of treatment on the outcome for the stratum that includes units whose potential post-treatment outcomes do not change (or change) under different treatment levels (Frangakis & Rubin, 2002; Mealli & Mattei, 2012).
@@
Most of the literature on principal stratification and principal causal effects has focused on the case of binary treatment and the binary post-treatment variable (see, e.g., Angrist et al., 1996). This implies that researchers deal with at most four principal strata that are the possible combinations of values of these two variables (see the broad literature for binary or discrete post-treatment variables, e.g., Cheng & Small, 2006; Imai, 2008; Ding et al., 2011; Mattei & Mealli, 2011; Mealli & Pacini, 2013; Mealli et al., 2016; Jiang et al., 2016, 2022; Mattei et al., 2024). Considering continuous post-treatment variables leads to a potentially infinite number of principal strata, which, in turn, makes inference and interpretation very challenging. Various solutions have been proposed in the causal inference literature. Proposals include: (i) dichotomize the post-treatment variable (Sjölander et al., 2009; Jiang et al., 2022), or model the continuous post-treatment variable conditional on the binary treatment using (ii) fully parametric or semiparametric modeling (Jin & Rubin, 2008; Conlon et al., 2014; Lu et al., 2023), or (iii) nonparametric modeling (Schwartz et al., 2011).
@@
Although these methods provide potential solutions to the challenges of continuous post-treatment variables, they often require strong assumptions or subjective choices. On the one hand, dichotomization of the post-treatment variable requires the choice of the threshold on the post-treatment variable (e.g., Zigler et al., 2012; Antonelli et al., 2023). Although this choice might be informed by the relevant subject matter literature, it entails a level of arbitrariness, and requires extensive checking of the robustness of the results with respect to the chosen thresholds.
Our proposed method introduces a novel and more interpretable approach to characterizing the principal strata, addressing some critical challenges in this framework. While, similarly Schwartz et al. (2011) modeled the continuous post-treatment variable using a Dirichlet process mixture (Ferguson, 1973), our approach advances this framework in two significant ways. First, we foster information sharing between the marginal distribution of potential post-treatment variables to enhance the ability to uncover the principal strata with a data-adaptive approach and provide a more interpretable definition of them. Second, while the Dirichlet process mixture proposed by Schwartz et al. (2011) adheres to the standard stick-breaking definition by Sethuraman (1994), our approach extends this by incorporating covariates into the weights of the mixture model. This adjustment significantly improves the imputation of missing potential post-treatment outcomes and enables strata identification without requiring predefined assumptions, consequently enhancing the estimation performance of treatment effects.
@@
To do so, we introduce a novel Bayesian approach to principal stratification. Specifically, we introduce a confounders-aware shared-atoms Bayesian mixture model endowed with innovative estimands for principal causal effects. Our model exploits the dependent Dirichlet process (Quintana et al., 2022) to establish a highly flexible distribution for the potential post-treatment outcomes given the confounders and the treatment variable, similar to the recent proposal of Zorzetto et al. (2024) in the context of heterogeneity in the treatment effects.
@@
Our proposed approach improves the current literature on nonparametric modeling for principal stratification in many respects. Firstly, it allows for the inclusion of the confounders in the weights of the infinite Bayesian mixture model, which provides a way to control for measured confounding and a more precise estimation of the causal effects. Secondly, the proposed model, which leverages shared random kernels’ atoms, allows us to identify associative/dissociative strata borrowing information across the treatment levels. Hence, our approach eliminates the need for subjective criteria to cluster principal strata. Instead, our model simultaneously (i) identifies the principal strata, according to the nature of the observed data, and (ii) imputes the missing post-treatment variables and outcomes. Moreover, our framework naturally incorporates uncertainty quantification for principal strata membership while providing robust estimation of principal causal effects.
This paper is organized as follows. In Section 2, we introduce the framework and the definition of the novel principal causal estimands. In Section 3, we propose our model and detail how it allows us to identify meaningful coarsening of the principal strata using a data-adaptive method. Theoretical prior properties provide insight into prior elicitation. An efficient Gibbs sampling algorithm to sample from the posterior distributions and estimate the model is also derived. Section 4 illustrates, via simulations, the ability of the proposed model to impute missing post-treatment outcomes, discover the main strata, and estimate the principal causal effects. In Section 5, we estimate the principal causal effects of air pollution policies on the mortality rate in the eastern United States (Zigler et al., 2018), where long-term fine particle exposure is the post-treatment variable. In Section 6, we conclude the paper with a discussion and avenues for future work.
2 Setup
@@
2.1 Identifying Assumptions
Following the principal stratification framework (Frangakis & Rubin, 2002; VanderWeele, 2011; Mealli & Mattei, 2012; Feller et al., 2017; Ding et al., 2017; Lu et al., 2023), the principal strata are not affected by treatment assignment; therefore, a principal stratification can be used as any unit classification to define meaningful causal estimates conditional on the principal strata and discover the heterogeneous treatment effect (Mealli & Mattei, 2012).
@@ 2.1 Identifying Assumptions
{assumption}
[Stable Unit Treatment Value Assumption]
@@
The Stable Unit Treatment Value Assumption (SUTVA) is a combination of (i) no interference between units, that is, the potential outcome and the potential values of the post-treatment variable of the unit $i$ do not depend on the treatment applied to other units, and (ii) consistency, that is, no different versions of the treatment levels assigned to each unit (Rubin, 1986). Under the principal stratification framework, this assumption is invoked for both the primary outcome variable and the post-treatment variable.
@@
{assumption}
[Strongly Ignorable Treatment Assignment] Given the observed covariates $\mathbf{x}_{i}$,
@@
$0<Pr\left(T_{i}=1\mid X_{i}=x\right)<1\>\>\forall\>x\in\mathcal{X},$
where $\mathcal{X}$ is the features’ space. This assumption states that: (i) for each unit, the potential primary outcomes and the potential post-treatment variables are independent of the treatment conditional on the set of covariates $X_{i}$; (ii) all units have a positive chance of receiving the treatment.
3 Confounders-Aware Shared-atoms Bayesian Mixture Model
@@
$\Pr(Y(0),Y(1),P(0),P(1),T,X)=\int_{\Theta}\prod_{i=1}^{n}\Pr(Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),T_{i},X_{i},\theta)p(\theta)d\theta,$
where $p(\theta)$ is the prior distribution for all the involved parameters $\theta$ that take values in the parameter space $\Theta$. The inner probability can be factorized as
$\Pr(T_{i}\mid Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),X_{i},\theta)$$\times\Pr(Y_{i}(0),Y_{i}(1)\mid P_{i}(0),P_{i}(1),X_{i},\theta)$
×$\times\Pr(P_{i}(0),P_{i}(1)\mid X_{i},\theta)$$\times\Pr(X_{i}\mid\theta).$ (1)
Invoking Assumption 2, the conditional probability of the treatment variable can be written as $\Pr(T_{i}\mid Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),X_{i},\theta)=\Pr(T_{i}\mid X_{i},\theta)$. Moreover, we replace the conditional distribution of the covariates with the empirical distribution so that $\Pr(X_{i}\mid\theta)=\Pr(X_{i})$.
@@
For modeling the post-treatment variable, we exploit a dependent nonparametric mixture, following the dependent Dirichlet process approach (MacEachern, 2000; Barrientos et al., 2012; Quintana et al., 2022; Denti et al., 2023). [^1] In particular, our model also shares some similarities with the hierarchical Dirichlet process (Teh et al., 2004, 2006; Teh & Jordan, 2009) and the recent common atom model of Denti et al. (2023). Specifically, we assume for each $i=1,\dots,n$:
@@
where ${\mathcal{K}}(\cdot;x,\psi)$ is a continuous density function, for every $\psi\in\Psi$, and $G_{\mathbf{x}_{i}}^{(t)}$ is a random probability measure depending on the confounders $x_{i}$ associated to an observation assigned to treatment level $t$. For each treatment level $t=\{0,1\}$, the distributions of $G_{\mathbf{x}_{i}}^{(0)}$ and $G_{\mathbf{x}_{i}}^{(1)}$ are characterized by specific weights assigned to a common set of atoms induced by the common process $\Pi$.
Following the characterization of the random probability measures, we write:
@@
$\pi_{l}^{(t)}(\mathbf{x}_{i})$ $=\omega_{l}^{(t)}(\mathbf{x}_{i})\prod_{r<l}\{1-\omega_{r}^{(t)}(\mathbf{x}_{i})\};$ (4)
@@
The discrete nature of the random probability measure $G_{\mathbf{x}_{i}}^{(t)}$, for $t=\{0,1\}$, allows us to introduce the latent categorical variables $S_{i}^{(t)}$, that describe the cluster allocations for each unit $i\in\{1,\dots,n\}$, whose clusters are defined by heterogeneous values for $P_{i}(t)$. Assuming $\mathds{P}\{S_{i}^{(t)}=l\}=\pi_{l}^{(t)}(\mathbf{x}_{i})$, we can write model in (2), exploiting conditioning on $S_{i}^{(t)}$, as
$\{P_{i}|\mathbf{x}_{i},t,\psi,S_{i}^{(t)}=l\}\sim{\mathcal{K}}(\cdot\mid\mathbf{x}_{i},\psi_{l}),\quad\psi_{l}\sim H.$
@@
Among the plethora of dependent non-parametric processes (see the recent review by Quintana et al., 2022, for a detailed description), we focused on the dependent probit stick-breaking process (Rodriguez & Dunson, 2011). Consistently with this, we specify:
@@
where $\Phi(\cdot)$ is the CDF of a standard Normal distribution, $\{\alpha_{l}^{(t)}(x_{i})\}_{l\geq 1}$ has Gaussian distributions with mean a linear combination of the $p$ covariates $x_{i}$, and $\{\beta_{l0}^{(t)},\beta_{l}^{(t)}\}_{l\geq 1}$ are the probit regression parameters treatment-specific.
We assume the kernel $\mathcal{K}$ to be Gaussian, so that model (2)–(3) becomes
$\{P_{i}(t)\mid\mathbf{x}_{i},S_{i}^{(t)}=l,\eta,\sigma\}\sim\mathcal{N}(\eta_{l},\sigma_{l}^{2}).$ (6)
@@
$\eta_{l}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(\mu_{\eta},\sigma_{\eta}^{2}),\mbox{ and }\sigma_{l}^{2}\stackrel{{\scriptstyle iid}}{{\sim}}\mbox{InvGamma}(\gamma_{1},\gamma_{2}).$
@@
For the regression parameters in (5), we assume the following multivariate Gaussian prior:
$\beta^{(t)}\sim\mathcal{N}_{(p+1)(L-1)}(\xi,\Omega)$ (7)
for $t=0,1$ and $l\geq 1$. According to Fasano et al. (2022), the Gaussian prior leads to a straightforward posterior computation, as discussed in the next section.

### v3

# Bayesian Nonparametrics for Principal Stratification with Continuous Post-Treatment Variables

@@ Abstract
Principal stratification provides a causal inference framework for investigating treatment effects in the presence of a post-treatment variable. Principal strata play a key role in characterizing the treatment effect by identifying groups of units with the same or similar values for the potential post-treatment variable at all treatment levels. The literature has focused mainly on binary post-treatment variables. Few papers considered continuous post-treatment variables. In the presence of a continuous post-treatment, a challenge is how to identify and characterize meaningful coarsening of the latent principal strata that lead to interpretable principal causal effects. This paper introduces the Confounders-Aware SHared atoms BAyesian mixture (CASBAH), a novel approach for principal stratification with binary treatment and continuous post-treatment variables. CASBAH leverages Bayesian nonparametric priors with an innovative hierarchical structure for the potential post-treatment outcomes that overcomes some of the limitations of previous works. Specifically, the novel features of our method allow for (i) identifying coarsened principal strata through a data-adaptive approach and (ii) providing a comprehensive quantification of the uncertainty surrounding stratum membership. Through Monte Carlo simulations, we show that the proposed methodology performs better than existing methods in characterizing the principal strata and estimating principal effects of the treatment. Finally, CASBAH is applied to a case study in which we estimate the causal effects of US national air quality regulations on pollution levels and health outcomes.
Bayesian Causal Inference, Principal Stratification, Shared Atom Mixture Model
@@ 1 Introduction
Principal stratification is a well-established causal inference approach that accounts for post-treatment variables (Frangakis and Rubin, 2002). A post-treatment variable is defined as a variable that is potentially affected by treatment, and it is also associated with the primary outcome. Thus, a post-treatment variable can be seen as an intermediate variable between the treatment and the primary outcome of interest. Post-treatment variables may pose inferential challenges for the estimation of treatment effects, but also provide a useful approach to conceptualizing the mechanisms/pathways through which the treatment affects the primary outcome through the intermediate variable.
@@
Principal stratification with respect to the post-treatment variable involves stratifying units into different groups, known as principal strata. Principal strata consist of units that share the same or similar potential values of the post-treatment variable under all possible treatment levels: by definition, principal strata are not affected by the treatment and thus represent an inherent latent characteristic of the units.
@@
Since the assignment of treatment does not influence the principal strata, comparisons of potential primary outcomes at different treatment levels within a principal stratum represent well-defined causal effects (Rubin, 1974). These specific comparisons are referred to as principal causal effects. Researchers have often referred to the dissociative (and associative) effect as the effect of treatment on the primary outcome for the stratum that includes units whose potential post-treatment variables do not change (or change) under different treatment levels (Frangakis and Rubin, 2002; Mealli and Mattei, 2012).
@@
Most of the literature on principal stratification and principal causal effects has focused on the case of binary treatment and the binary post-treatment variable (see, e.g., Angrist et al., 1996). This implies that researchers deal with at most four principal strata that are the possible combinations of values of these two variables (see the broad literature for binary or discrete post-treatment variables, e.g., Cheng and Small, 2006; Imai, 2008; Ding et al., 2011; Mattei and Mealli, 2011; Mealli and Pacini, 2013; Mealli et al., 2016; Jiang et al., 2016, 2022; Mattei et al., 2024). Considering continuous post-treatment variables leads to a potentially infinite number of principal strata, which, in turn, makes inference and interpretation very challenging. Various solutions have been proposed in the causal inference literature. Proposals include: (i) dichotomize the post-treatment variable (Sjölander et al., 2009; Jiang et al., 2022), or model the continuous post-treatment variable conditional on the binary treatment using (ii) fully parametric or semiparametric modeling (Jin and Rubin, 2008; Conlon et al., 2014; Lu et al., 2023; Schwartz et al., 2011), or (iii) nonparametric modeling (Antonelli et al., 2023).
@@
Although these methods provide potential solutions to the challenges of continuous post-treatment variables, they often require strong assumptions or subjective choices. For example, dichotomization of the post-treatment variable requires the choice of the threshold on the post-treatment variable (e.g., Zigler et al., 2012; Antonelli et al., 2023). Although this choice might be informed by the relevant subject matter literature, it entails a level of arbitrariness/subjectivity and requires extensive checking of the robustness of the results with respect to the chosen thresholds.
@@
Our proposed method introduces a novel and more interpretable approach to characterizing the principal strata, addressing some critical challenges in this framework. To do so, we propose a novel Bayesian approach to principal stratification. Specifically, we introduce a Confounders-Aware SHared-atoms BAyesian mixture model (denoted by the acronym CASBAH, with the intentional alteration of the letter “H") endowed with innovative estimands for principal causal effects. CASBAH exploits the dependent Dirichlet process (MacEachern, 1999; Quintana et al., 2022) to establish a highly flexible distribution for the potential post-treatment variable given the confounders and the treatment variable, similar to the recent proposal of Zorzetto et al. (2024) in the context of heterogeneity in the treatment effects.
@@
Our proposed approach improves the current literature on nonparametric modeling for principal stratification in many respects. First, it allows for the inclusion of the confounders in the weights of the infinite Bayesian mixture model, which provides a way to control for measured confounding and a more precise estimation of the causal effects. Secondly, CASBAH, which leverages shared random kernels’ atoms, allows us to identify coarsened associative/dissociative strata borrowing information across the treatment levels. Hence, our approach eliminates the need for subjective criteria to cluster principal strata. Instead, CASBAH simultaneously (i) identifies the principal strata, according to the nature of the observed data, and (ii) imputes the missing post-treatment variables and primary outcomes. Moreover, our framework naturally incorporates uncertainty quantification for principal strata membership while providing robust estimation of principal causal effects.
2 Causal Inference Setup
@@
Following the principal stratification framework (Frangakis and Rubin, 2002; VanderWeele, 2011; Mealli and Mattei, 2012; Feller et al., 2017; Ding et al., 2017; Lu et al., 2023), the principal strata are not affected by treatment assignment; therefore, a principal stratification can be used as any unit classification to define meaningful causal estimates conditional on the principal strata and discover the heterogeneous treatment effect (Mealli and Mattei, 2012).
@@
Assumption 1 (Stable Unit Treatment Value Assumption)
@@
The Stable Unit Treatment Value Assumption is a combination of (i) no interference between units, that is, the potential outcome and the potential values of the post-treatment variable of the unit $i$ do not depend on the treatment applied to other units, and (ii) consistency, that is, no different versions of the treatment levels assigned to each unit (Rubin, 1986). Under the principal stratification framework, this assumption is invoked for both the primary outcome variable and the post-treatment variable.
@@
Assumption 2 (Strongly Ignorable Treatment Assignment)
Given the observed covariates $\mathbf{x}_{i}$,
@@
$0<\text{P}\left(T_{i}=1\mid X_{i}=x\right)<1\>\>\forall\>x\in\mathcal{X}.$
Assumption 2 states that: (i) for each unit, the potential primary outcomes and the potential post-treatment variables are independent of the treatment conditional on the set of covariates $X_{i}$; (ii) all units have a positive chance of receiving the treatment.
3 Model
3.1 Confounders-Aware Shared-atoms Bayesian Mixture Model
@@
$\text{P}(Y(0),Y(1),P(0),P(1),T,X)=\int_{\Theta}\prod_{i=1}^{n}\text{P}(Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),T_{i},X_{i},\theta)\text{P}(\theta)d\theta,$
where $\text{P}(\theta)$ is the prior measure for all the involved parameters $\theta$ that take values in the parameter space $\Theta$. The inner probability can be factorized as
$\text{P}(T_{i}\mid Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),X_{i},\theta)$ $\times\text{P}(Y_{i}(0),Y_{i}(1)\mid P_{i}(0),P_{i}(1),X_{i},\theta)$
$\times\text{P}(P_{i}(0),P_{i}(1)\mid X_{i},\theta)$ $\times\text{P}(X_{i}\mid\theta).$ (1)
Invoking Assumption 2, the conditional probability of the treatment variable can be written as $\text{P}(T_{i}\mid Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),X_{i},\theta)=\text{P}(T_{i}\mid X_{i},\theta)$. Moreover, we replace the conditional distribution of the covariates with the empirical distribution so that $\text{P}(X_{i}\mid\theta)=\text{P}(X_{i})$.
@@
For modeling the post-treatment variable, we exploit a dependent nonparametric mixture, following the dependent Dirichlet process approach (MacEachern, 2000; Barrientos et al., 2012; Quintana et al., 2022). In particular, our model also shares some similarities with the hierarchical Dirichlet process (Teh et al., 2004, 2006; Teh and Jordan, 2009) and the recent common atom model of Denti et al. (2023). Specifically, we assume for each $i=1,\dots,n$:
@@
where ${\mathcal{K}}(\cdot;\psi)$ is a continuous density function, for every $\psi\in\Psi$, and $G_{\mathbf{x}_{i}}^{(t)}$ is a random probability measure depending on the confounders $x_{i}$ associated to an observation assigned to treatment level $t$. The random probability measures $G_{\mathbf{x}_{i}}^{(0)}$ and $G_{\mathbf{x}_{i}}^{(1)}$ are defined as
@@
$\pi_{l}^{(t)}(\mathbf{x}_{i})$ $=\omega_{l}^{(t)}(\mathbf{x}_{i})\prod_{r<l}\{1-\omega_{r}^{(t)}(\mathbf{x}_{i})\},$ (4)
@@
The discrete nature of the random probability measure $G_{\mathbf{x}_{i}}^{(t)}$, for $t=\{0,1\}$, allows us to introduce the latent categorical variables $S_{i}^{(t)}$, that describe the cluster allocations for each unit $i\in\{1,\dots,n\}$, whose clusters are defined by heterogeneous values for $P_{i}(t)$. Assuming $\Pr(S_{i}^{(t)}=l)=\pi_{l}^{(t)}(\mathbf{x}_{i})$, we can write model in (2), exploiting conditioning on $S_{i}^{(t)}$, as
$\{P_{i}|\mathbf{x}_{i},t,\psi,S_{i}^{(t)}=l\}\sim{\mathcal{K}}(\cdot\mid\mathbf{x}_{i},\psi_{l}),\qquad\psi_{l}\sim H.$
@@
Among the plethora of dependent non-parametric processes, we focus on the probit stick-breaking process (Rodriguez and Dunson, 2011) and specifically assuming
@@
where $\Phi(\cdot)$ is the cumulative distribution function of a standard Gaussian distribution, $\{\alpha_{l}^{(t)}(x_{i})\}_{l\geq 1}$ has Gaussian distributions with mean a linear combination of the $p$ covariates $x_{i}$, and $\{\beta_{l0}^{(t)},\beta_{l}^{(t)}\}_{l\geq 1}$ are the treatment-specific probit regression parameters. For the regression parameters in (5), we assume the following multivariate Gaussian prior:
$\beta^{(t)}\sim\mathcal{N}_{(p+1)(L-1)}(\xi,\Omega),$ (6)
for $t=0,1$ and $l\geq 1$. Consistently with Fasano et al. (2022), the Gaussian prior leads to a straightforward posterior computation, as discussed in the next section.
We assume the kernel density to be Gaussian, so that model (2)–(3) becomes
$\{P_{i}(t)\mid S_{i}^{(t)}=l,\eta,\sigma\}\sim\mathcal{N}(\eta_{l},\sigma_{l}^{2}).$ (7)
@@
$\eta_{l}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(\mu_{\eta},\sigma_{\eta}^{2}),\mbox{ and }\sigma_{l}^{2}\stackrel{{\scriptstyle iid}}{{\sim}}\mbox{InvGamma}(\gamma_{1},\gamma_{2}),$